Mr. Castro gave his daughter 300 pesos on her 10th birthday, and intends to increase this by 200 pesos each year. How much will he daughter receive on her 18th birthday
Answer by checkley77(12844)

300+(8*200)=300+1,600=1,900 FOR THE 20 TH. BIRTHDAY